# On the Need to Rethink Trust in AI Assistants for Software Development: A Critical Review

### v2

@@ Abstract
Trust is a fundamental concept in human decision-making and collaboration that has long been studied in philosophy and psychology. However, software engineering (SE) articles often use the term ‘trust’ informally—providing an explicit definition or embedding results in established trust models is rare. In SE research on AI assistants, this practice culminates in equating trust with the likelihood of accepting generated content, which, in isolation, does not capture the full complexity of the trust concept. Without a common definition, true secondary research on trust is impossible. The objectives of our research were: (1) to present the psychological and philosophical foundations of human trust, (2) to systematically study how trust is conceptualized in SE and the related disciplines human-computer interaction and information systems, and (3) to discuss limitations of equating trust with content acceptance, outlining how SE research can adopt existing trust models to overcome the widespread informal use of the term ‘trust’. We conducted a literature review across disciplines and a critical review of recent SE articles focusing on conceptualizations of trust. We found that trust is rarely defined or conceptualized in SE articles. Related disciplines commonly embed their methodology and results in established trust models, clearly distinguishing, for example, between initial trust and trust formation and between appropriate and inappropriate trust. On a meta-scientific level, other disciplines further discuss whether and when trust can be applied to AI assistants at all. Our study reveals a significant maturity gap of trust research in SE compared to related disciplines. We provide concrete recommendations on how SE researchers can adopt established trust models and instruments to study trust in AI assistants beyond the acceptance of generated software artifacts.
Index Terms:
Software Engineering, AI Assistants, Large Language Models, Human-Computer Interaction, Information Systems, Trust, Literature Review, Critical Review
@@ I Introduction
Figure 1: Trust model by [95] applied to AI assistants (cf. Figure 2) with references to existing measurement instruments.
@@
Since the launch of ChatGPT in November 2022 and the subsequent hype around generative artificial intelligence (GenAI), AI assistants that help humans generate images, text, or source code have become an integral part of the daily work of many knowledge workers around the world [97]. GenAI tools including ChatGPT and GitHub Copilot support software engineers in tasks such as understanding existing code, fixing bugs, or implementing new features [140, 74]. GitHub claims that, as of February 2023, Copilot is behind an average of 46 percent of developers’ code across all programming languages. [^1] GitHub Blog: GitHub Copilot now has a better AI model The IDE integration of GitHub Copilot and similar AI development assistants allows tool vendors to collect fine-grained usage data to understand which suggestions developers accept, which they reject, and which they modify before committing a change [169].
@@
Previous work coauthored by employees of companies with commercial interests in increasing tool adoption, including Microsoft, Google, and Amazon, investigated what they refer to as trust in AI assistants [31, 154, 24, 104, 120, 66]. In some of those articles, the concept of trust is operationalized as tool adoption in general or, more specifically, as a high acceptance rate for generated content. In short, they equate trust with a high probability of developers accepting generated software artifacts, while ignoring trust formation and the antecedents of trust (see Figure 1). Compared to existing conceptualizations of trust in foundational disciplines such as psychology and philosophy, but also more applied disciplines such as human-computer interaction (HCI) and information systems (IS), we consider this view to be too simplistic. It disregards decades of research on trust.
@@
Although embedding software engineering (SE) research in existing models and theories of trust is—as our literature review will show—not widespread, our community does have a history of discussing trust in source code. For example, in his Turing Award lecture in 1984, Ken Thompson addressed the morality of trust, stating: You can’t trust code that you did not totally create yourself… No amount of source-level verification or scrutiny will protect you from using untrusted code [147]. Given the widespread adoption of AI-based code generation and the persistent issue of hallucinations [14], it becomes imperative to adopt a more holistic perspective on trust, one that extends beyond the notion of accepting AI-generated content. Combining Ken Thompson’s statement with the increasing amount of generated code, this leads to a critical question: Can developers genuinely trust generated software artifacts or is there a conceptual difference between trust and mere acceptance of generated content?
@@
Our analysis of conceptualizations of trust in other disciplines reveals that viewing trust as mere acceptance of generated content is problematic for two key reasons. First, it lacks a solid theoretical foundation because it is unclear whether what is being measured is actually aligned with common conceptualizations of trust. People’s behavior, and hence their likelihood of accepting generated content, can be influenced not only by the perceived trustworthiness of the AI they are using, but also by other factors (see Figure 1 for an example). Factors that affect trust are not necessarily directly related to trust [80, 95]. Such factors include people’s specific situation (e.g., time pressure), the type of AI system they are interacting with, the specific task they are performing, and the potential consequences of decisions. Second, viewing trust as mere acceptance of generated content is potentially harmful. Even if trust could be measured through acceptance alone, this would not imply that the measured trust is appropriate trust. Given the potential maintainability challenges posed by generated code [87], security risks associated with misplaced trust in AI-generated code [111], and the potential negative impact on developers’ learning process [162, 131], research needs to distinguish appropriate from inappropriate trust.
@@
We advocate for a broader interdisciplinary approach to understanding trust in SE. Note that, in this paper, we focus on human/human-like trust, not trust in a more technical sense (e.g., chains of trust in a security context). Our goal is to motivate the need of our community to rethink how we conceptualize and measure trust, especially in the context of AI assistants for software development. Our target audience are SE researchers planning to design studies on trust as well as SE editors and reviewers who are responsible for ensuring our communities’ epistemological and methodological standards when it comes to researching interdisciplinary topics such as trust. To start a community discourse on rethinking trust and to support researchers who want to adopt and adapt established trust theories, we address the following research questions:
@@
After introducing the foundations of human trust based on research in psychology and philosophy (Section II), we present the results of a literature review on conceptualizations of trust (Section III-B) in HCI and IS (RQ1) as well as SE (RQ2). We complement that literature review with a critical review of trust conceptualizations (Section III-C) in recent SE articles (RQ2) and then synthesize the results of those two literature reviews into a discussion on how to rethink trust in AI assistants for software development (RQ3), providing actionable recommendations (Sections IV and VI).
@@ II Background: Foundations of Trust
Trust and the related concept of trustworthiness have been discussed and conceptualized long before disciplines such as SE, HCI, or IS recognized these concepts. Therefore, before we dive into our cross-disciplinary literature review of trust, we outline how philosophy and psychology, that is, disciplines which have been studying the foundations of human behavior for a long time, conceptualize human trust.
First, it is important to distinguish trust from trustworthiness. Both philosophy and psychology agree that trust is an attitude that one entity (the trustor) has towards another entity (the trustee) that it hopes will be trustworthy. Hence, leaving perceived trustworthiness aside, trustworthiness is considered a property of the trustee and not an attitude [101, 80, 102]. Furthermore, both disciplines agree that trust is a contextual matter: a trustor may trust a trustee in one situation or for one specific task while, at the same time, not trusting them in another situation or for another task. Typically, the trustor is human, while it is debatable whether the trustee needs to be human as well (see Section II-B for details).
@@ II-A Psychology
Figure 2: Trust model by [95] (own illustration based on [95]).
@@
In this model, ability refers to the trustee’s competence in a given domain, benevolence captures the trustee’s (assumed) goodwill and positive intentions towards the trustor, and integrity refers to the trustee’s adherence to ethical principles the trustor agrees with. The model suggests that trust is high when all three components of trustworthiness are perceived as high by the trustor, whereas deficiencies in any factor can weaken trust. How much (changes in) the factors affect trust depends on the trustor’s propensity to trust. This propensity also influences the trustor’s initial level of trust. Trust can manifest itself in specific situations, depending on the level of trust and the risk perceived by a trustor when they trust the trustee to do a certain task. Depending on the outcome of this risk-taking, the perception of the three factors can change.
Figure 3: Trust model by [80] (see Figure 2 in [80]).
[80] take this model, expand it, and apply it to automation contexts [80]. They introduce the concepts of calibration, resolution, and specificity (see Figure 3). Calibration concerns the crucial question of whether the trust a trustor has corresponds to the real capabilities, that is, the actual trustworthiness of a system. Ill-calibrated trust may lead to disuse (in the case of dis-/undertrust) or misuse (in the case of overtrust) of a system. Resolution is about how changes in system capabilities change levels of trust. For example, with low resolution, high changes in capability would lead to small changes in trust. Finally, specificity reflects how granular the trust is. Specificity emphasizes the context-dependence of trust by stating how much trust depends on the capabilities of a particular subsystem or the situation. According to [80], good calibration, high resolution, and high specificity of trust can mitigate misuse and disuse of automation systems and enhance human-automation partnerships. Although they also introduce a model on the interaction of context, agent characteristics, and cognitive properties with the appropriateness of trust, their work is mostly cited for making the above distinctions.
@@ II-B Philosophy
[101] has analyzed philosophical debates on trust and distills three requirements for trust [101]. In summary, trust requires that one can:
@@
For most philosophers, trust is a kind of reliance, although it involves some extra factor [101]. This extra factor is often determined by how philosophers interpret the three requirements listed above.
For example, regarding the factor of vulnerability to betrayal, trust is often seen as pre-supposing anti-monitoring (see, e.g., [10, 73]). The idea here is that if a person monitors an entity, this person cannot be betrayed by that entity. At most, the person could be disappointed if the entity does not perform as intended. Here, there is an interesting difference to [95]’s theory of trust, which explicitly decouples monitoring from trust. However, their theory is flexible enough to model monitoring situations. For example, an employee may devalue the benevolence of (and thus the trust in) their employer due to constant monitoring.
The factors of competence and willingness (i.e., motivation) have received the most attention in the debate, with many competing theories that exist [101]. These theories range from the trustee acting out of self-interest [53] or goodwill [68, 10] to theories of normative expectation [44]. In summary, while there is a minimum consensus of the three requirements mentioned above, philosophers differ so much in the exact formulations of these requirements that some even assume that there are different types of trust. This would mean that trust is not one form of reliance but different ones [136, 124, 30].
Philosophers place more emphasis on the preconditions of trust or the characteristics of its formation than psychologists. Therefore, in addition to the concept of calibrated trust (which they call well-grounded trust), they also introduce the concepts of justified and plausible trust [101]. Justification is about the indicators that have been consulted for trust formation, that is, whether it is based on good evidence. If a person happens to trust the system in the right way, but does so by chance without any indications, then philosophers would not speak of justified trust. Finally, plausibility is about whether the preconditions for trust are satisfied. For example, if two parties are antagonistic, there can be no trust between them.
@@ II-C Philosophy and Psychology on Trust and AI
Philosophical and psychological discussions about trust continue to this day, often in specific areas of application. AI is one of them, especially in recent times. There are two discussions that are worth highlighting here. In particular, with regard to the plausibility of trust, there are discussions as to whether it makes sense to apply the term trust to AI at all. From a philosophical point of view, AI has no motivation with which it could betray us.
@@
This discussion even made it into the media when the Guidelines for Trustworthy AI of the high-level AI expert group set up by the European Commission were published. One of the members of this group, the German philosopher Thomas Metzinger, gave an interview shortly before the guidelines were published, in which he complained about the strong involvement of companies in the design of the guidelines. He claimed that these would use the principle of trustworthy AI for ethics washing, because in his opinion, the term trustworthy AI is conceptual nonsense. [^2] EU guidelines: Ethics washing made in Europe (Tagesspiegel) While some authors agree with Metzinger in that the concept of trust cannot be applied to AI systems, for example, since they do not possess emotive states or cannot be held responsible [119, 37], others defend the possibility to trust AI, for instance, by showing how philosophical theories of trust can be adapted to AI [151, 121, 62, 30, 107] or by differentiating trust in AI from trust in humans [166] (see [38] for an overview of the debate).
@@
Irrespective of the discussion as to whether the concept of trust can be meaningfully applied to AI, the Guidelines for Trustworthy AI had strong effects on the subsequent operationalization of trust in AI. By specifying requirements that are essential for trustworthiness, they provide fertile ground for psychological and philosophical research to operationalize these requirements more precisely (see, e.g., [93, 84]).
@@
The other noteworthy discussion revolves around the relationship between explainable AI (XAI) and trust in AI. Increasing trust is one of—if not the–primary goal of XAI [70, 79]. [70] refer to this as the explainability-trust hypothesis [70]. However, studies show that providing explanations does not consistently increase trust, and in many cases, even decreases it [70, 79]. Here, trust calibration plays an important role, as explanations that reveal system errors should lead to lower trust. Unfortunately, psychological biases influence this process, as people sometimes trust a system simply because of the illusion of an explanation [40]. On the same topic, the discussion about the plausibility of trust on the philosophical side is worth noting. For example, [41] argue, drawing on the concept of anti-monitoring, that explainability cannot directly contribute to trust, as it constitutes a form of monitoring [41].
@@
Regarding biases and the possibility of generating problematic content, there is now more and more research that argues that, for a system to be adequately trusted, it is not just the system that needs to function correctly. Rather, for a system to be perceived as benevolent and integer, there must be regulation, auditing processes, and established standards that govern the entire engineering process [157, 91, 112]. [157], for instance, argue that ethical governance, consisting of ethical codes of conduct, ethics training, responsible innovation, and transparency, is essential to build trust in AI systems [157]. Similarly, [91] argue that user trust requires evidence on the functionality of an AI assistant at the level of its design, the practices of the developing organization, and oversight by external bodies [91].

### v3

@@ Abstract
Trust is a fundamental concept in human decision-making and collaboration that has long been studied in philosophy and psychology. However, software engineering (SE) articles often use the term trust informally; providing an explicit definition or embedding results in established trust models is rare. In SE research on AI assistants, this practice culminates in equating trust with the likelihood of accepting generated content, which, in isolation, does not capture the full conceptual complexity of trust. Without a common definition, true secondary research on trust is impossible. The objectives of our research were: (1) to present the psychological and philosophical foundations of human trust, (2) to systematically study how trust is conceptualized in SE and the related disciplines human-computer interaction and information systems, and (3) to discuss limitations of equating trust with content acceptance, outlining how SE research can adopt existing trust models to overcome the widespread informal use of the term trust. We conducted a literature review across disciplines and a critical review of recent SE articles with a focus on trust conceptualizations. We found that trust is rarely defined or conceptualized in SE articles. Related disciplines commonly embed their methodology and results in established trust models, clearly distinguishing, for example, between initial trust and trust formation and between appropriate and inappropriate trust. On a meta-scientific level, other disciplines even discuss whether and when trust can be applied to AI assistants at all. Our study reveals a significant maturity gap of trust research in SE compared to other disciplines. We provide concrete recommendations on how SE researchers can adopt established trust models and instruments to study trust in AI assistants beyond the acceptance of generated software artifacts.
@@ I Introduction
Figure 1: Trust model by [Mayer1995] applied to AI assistants with references to existing measurement instruments [MayerDavies1999, DBLP:journals/isr/McKnightCK02, Frazier01102013, Scholz17012025, DavidSchoorman02012016, gillespie201520, DBLP:journals/jds/SchmidtBT20].
@@
Since the launch of ChatGPT in November 2022 and the subsequent hype around generative artificial intelligence (GenAI), AI assistants that help humans generate images, text, or source code have become an integral part of the daily work of many knowledge workers around the world [mcgeorge2023chatgpt]. GenAI tools, for example ChatGPT and GitHub Copilot, support software engineers in tasks such as understanding existing code, fixing bugs, or implementing new features [sridhara2023chatgpt, khojah2024beyond]. GitHub claims that, as of February 2023, Copilot is behind an average of 46 percent of developers’ code across all programming languages. [^1] GitHub Blog: GitHub Copilot now has a better AI model The IDE integration of GitHub Copilot and similar AI development assistants allows tool vendors to collect fine-grained usage data to understand which suggestions developers accept, which they reject, and which they modify before committing a change [DBLP:journals/cacm/ZieglerKLRRSSA24].
@@
Previous work coauthored by employees of companies with commercial interests in increasing tool adoption, including Microsoft, Google, and Amazon, investigated what they refer to as trust in AI assistants [cheng2024would, 12, brown2024identifying, DBLP:journals/software/MurilloEDBKMM24, Sabouri2025, DBLP:conf/icse/JohnsonBFFZ23]. In some of those articles, the concept of trust is operationalized as tool adoption in general or, more specifically, as a high acceptance rate for generated content. In short, they equate trust with a high probability of developers accepting generated software artifacts, while ignoring aspects like trust formation and the antecedents of trust (see Figure 1). Compared to existing conceptualizations of trust in foundational disciplines such as psychology and philosophy, but also more applied disciplines such as human-computer interaction (HCI) and information systems (IS), we consider this view to be too simplistic. It disregards decades of research on trust.
@@
Although embedding software engineering (SE) research in existing models and theories of trust is not widespread (see Section III), our community does have a history of discussing trust in source code. For example, in his Turing Award lecture in 1984, Ken Thompson addressed trust in foreign code, stating: You can’t trust code that you did not totally create yourself… No amount of source-level verification or scrutiny will protect you from using untrusted code [DBLP:journals/cacm/Thompson84]. Given the widespread adoption of AI-based code generation and the persistent issue of hallucinations [DBLP:journals/corr/abs-2409-05746], it becomes imperative to adopt a more holistic perspective on trust, one that extends beyond the notion of accepting AI-generated content. Combining Ken Thompson’s statement with the increasing amount of generated code, this leads to a critical question: Can developers genuinely trust generated software artifacts, or is there a conceptual difference between trust and mere acceptance of generated content?
@@
Our analysis of conceptualizations of trust in other disciplines reveals that viewing trust as mere acceptance of generated content is problematic for two key reasons. First, it lacks a solid theoretical foundation because it is unclear whether what is being measured is actually aligned with common conceptualizations of trust. People’s behavior, and hence their likelihood of accepting generated content, can be influenced not only by the perceived trustworthiness of the AI they are using, but also by other factors (see Figure 1 for an example). Factors that affect trust are not necessarily directly related to trust [Lee2004Trust, Mayer1995]. Such factors include people’s specific situation (e.g., time pressure), the type of AI system with which they are interacting, the specific task they are performing, and the potential consequences of decisions.
Second, viewing trust as mere acceptance of generated content is potentially harmful. Even if trust could be measured through acceptance alone, this would not imply that the measured trust is appropriate trust. Given the potential maintainability challenges posed by generated code [DBLP:journals/tosem/LiuLWTLLL24], security risks associated with misplaced trust in AI-generated code [DBLP:conf/ccs/PerryS0B23, DBLP:conf/sp/PearceA0DK22, DBLP:conf/uss/SandovalPNKGD23, DBLP:conf/chi/SerafiniYN25, DBLP:conf/sac/KudriavtsevaHG25, DBLP:conf/wcre/MajdinasabBRDTK24, DBLP:conf/ccs/KlemmerHPLBPMRV24, DBLP:conf/sp/OhLPKK24], and the potential negative impact on developers’ learning process [14, DBLP:journals/cacm/Shein24b], researcher need to distinguish appropriate from inappropriate trust. The fact that many developers distrust the accuracy of AI assistants2222025 Stack Overflow Developer Survey but nevertheless adopt these tools indicates that there is more to adoption than just trusting a tool.
We advocate for a broader interdisciplinary approach to understanding trust in SE. Our goal is to motivate the SE research community to rethink how we conceptualize and measure trust, especially in the context of AI assistants for software development. We would like to establish that software cannot be developed with the goal of “increasing trust” without knowing what trust actually entails. Without a scientifically validated trust model, trust becomes an empty marketing concept or, even worse, a concept used for ethics washing (see Section II-C). Without an established trust model, there is no reference point to compare claims of “increased trust” to and check whether they are actually true. More specifically, without a clear understanding of the concept of trust, it is difficult to consider trust in design goals, which in turn makes it difficult to design interactive AI assistants for software development in a way that addresses concrete trust-related challenges.
@@
Note that in this paper, we focus on human or human-like trust, not trust in a more technical sense (e.g., the implementation of chains of trust in a security context). Our target audience are SE researchers planning to design studies on trust as well as editors and reviewers of SE research who are responsible for ensuring our communities’ epistemological and methodological standards when it comes to researching interdisciplinary topics such as trust. To start a community discourse on rethinking trust and to support researchers who want to adopt and adapt established trust models, we address the following research questions (RQs):
@@
After introducing the foundations of human trust based on research in psychology and philosophy (Section II), we present the results of a literature review (Section III) on conceptualizations of trust in HCI and IS (RQ1) as well as SE (RQ2). We complement that literature review with a critical review of trust conceptualizations in recent SE articles (RQ2) and then synthesize the results of the two literature reviews into a discussion on how to rethink trust in AI assistants for software development (RQ3), including actionable recommendations for SE researchers (Section IV). We then discuss threats to validity (Section V) and conclude the paper (Section VI).
@@ II Background: Foundations of Trust
Trust and the related concept of trustworthiness have been discussed and conceptualized long before disciplines such as SE, HCI, or IS recognized these concepts. Therefore, before we dive into our cross-disciplinary literature review of trust, we outline how philosophy and psychology, that is, disciplines that have been studying the foundations of human behavior for a long time, conceptualize human trust.
First, it is important to distinguish trust from trustworthiness. Both philosophy and psychology agree that trust is an attitude that one entity (the trustor) has toward another entity (the trustee) that it hopes will be trustworthy. Hence, leaving perceived trustworthiness aside, trustworthiness is considered a property of the trustee and not an attitude [McLeod2023Trust, Lee2004Trust, 10.1145/3696449]. Furthermore, both disciplines agree that trust and trustworthiness are contextual matters: a trustor may trust a trustee in one situation or for one specific task while, at the same time, not trusting them in another situation or for another task. Typically, the trustor is human, while it is debatable whether the trustee must also be human (see Section II-B).
@@ II-A Psychology
Figure 2: Trust model by [Mayer1995] (own illustration based on [Mayer1995]).
@@
In this model, ability refers to the trustee’s competence in a given domain, benevolence captures the trustee’s (assumed) goodwill and positive intentions toward the trustor, and integrity refers to the trustee’s adherence to ethical principles the trustor agrees with. The model suggests that trust is high when all these three components of trustworthiness are perceived by the trustor as high, while deficiencies in any factor can weaken trust. How much (changes in) the factors affect trust depends on the trustor’s propensity to trust. This propensity also influences the trustor’s initial level of trust. Trust can manifest itself in specific situations, depending on the level of trust and the risk perceived by a trustor when they trust the trustee to perform a certain task. Depending on the outcome of this risk-taking, the perception of the three factors can change.
TrustAutomation Capability(Trustworthiness) Overtrust: Trust exceeds system capabilities, leading to misuse Calibrated trust: Trust matches system capabilities, leading to appropriate use Distrust: Trust falls short of system capabilities, leading to disuse Good resolution: A range of system capability maps onto the same range of trust Poor resolution: A large range of system capability maps onto a small range of trust
Figure 3: Trust model by [Lee2004Trust] (own illustration based on [Lee2004Trust]).
[Lee2004Trust] take this model, expand it and apply it to automation contexts [Lee2004Trust]. They introduce the concepts of calibration, resolution, and specificity (see Figure 3). Calibration concerns the crucial question of whether the trust a trustor has corresponds to the real capabilities, that is, the actual trustworthiness of a system. Ill-calibrated trust may lead to disuse (in the case of dis-/undertrust) or misuse (in the case of overtrust) of a system. Resolution is about how changes in system capabilities change levels of trust. For example, with low resolution, high changes in capability would lead to small changes in trust. Finally, specificity reflects how granular the trust is. Specificity emphasizes the context-dependence of trust by stating how much trust depends on the capabilities of a particular system or the situation. According to [Lee2004Trust], good calibration, high resolution, and high specificity of trust can mitigate misuse and disuse of automation systems and enhance human-automation partnerships. Although they also introduce a model on the interaction of context, agent characteristics, and cognitive properties with the appropriateness of trust, their work is mostly cited for making the above distinctions.
@@ II-B Philosophy
[McLeod2023Trust] has analyzed philosophical debates on trust and distills three requirements for trust [McLeod2023Trust]. In summary, trust requires that one can:
@@
For most philosophers, trust is a kind of reliance, although it involves some extra factor [McLeod2023Trust]. This extra factor is often determined by how philosophers interpret the three requirements listed above.
For example, regarding the factor of vulnerability to betrayal, trust is often connected to anti-monitoring (see, e.g., [Baier1986Trust, Keren2014Trust]). The idea here is that if a person monitors an entity, this person cannot be betrayed by that entity. At most, the person could be disappointed if the entity does not perform as intended. Here, there is an interesting difference to [Mayer1995]’s theory of trust, which explicitly decouples monitoring from trust. However, their theory is flexible enough to capture the same intuition concerning monitoring situations. For example, an employee may devalue the benevolence of (and thus the trust in) their employer due to constant monitoring.
The factors of competence and willingness (i.e., motivation) have received the most attention in the debate, with many competing theories that exist [McLeod2023Trust]. These theories range from the trustee acting out of self-interest [Hardin2002Trust] or goodwill [Jones1999Second-Hand, Baier1986Trust] to theories of networks of trust [Frost2014Imposters]. In summary, while there is a minimum consensus of the three requirements mentioned above, philosophers differ so much in the exact formulations of these requirements that some even assume that there are different types of trust. This would mean that trust is not one form of reliance but several different ones [Simpson2012Trust, Scheman2020Trust, 1].
Philosophers place more emphasis on the preconditions of trust and the conditions of its formation than psychologists. Therefore, in addition to the concept of calibrated trust (which they call well-grounded trust), they also introduce the concepts of justified and plausible trust [McLeod2023Trust]. Justification is about the indicators that have been consulted for trust formation, that is, whether it is based on good evidence. If a person happens to trust the system in the right way, but does so by chance without any indications, then philosophers would not speak of justified trust. Finally, plausibility is about whether the preconditions for trust are satisfied. For example, if two parties are antagonistic, there can be no trust between them.
@@ II-C Philosophy and Psychology on Trust and AI
Philosophical and psychological discussions about trust continue to this day, often in specific areas of application. AI is one of them, especially in recent times. There are three discussions that are worth highlighting here. In particular, with regard to the plausibility of trust, there are discussions as to whether it makes sense to apply the term trust to AI at all. From a philosophical perspective, current AI is not conscious, so it does not make sense to speak of it ‘betraying’ us.
@@
This discussion even made it into the media when the Guidelines for Trustworthy AI of the high-level AI expert group set up by the European Commission were published. One of the members of this group, the German philosopher Thomas Metzinger, gave an interview shortly before the guidelines were published, in which he complained about the strong involvement of companies in the design of the guidelines. He claimed that these would use the principle of trustworthy AI for ethics washing, because in his opinion, the term trustworthy AI is conceptual nonsense. [^3] EU guidelines: Ethics washing made in Europe (Tagesspiegel) While some authors agree with Metzinger in that the concept of trust cannot be applied to AI systems, for example, since they do not possess emotive states or cannot be held responsible [10, Dorsch2024Quasi], others defend the possibility to trust AI, for instance, by showing how philosophical theories of trust can be adapted to AI [Viehoff2023Making, Safdari2024Toward, 4, 1, 7] or by differentiating trust in AI from trust in humans [Zanotti2023Keep] (see [2] for an overview of the debate).
@@
This leads to a second, more practice-oriented discussion: How is trustworthiness operationalized? Irrespective of the discussion as to whether the concept of trust can be meaningfully applied to AI, the Guidelines for Trustworthy AI had strong effects on the subsequent operationalization of AI trustworthiness (and with that trust in AI). By specifying requirements that are essential for trustworthiness, e.g. technical robustness, safety, transparency, and accountability, they provide fertile ground for psychological and philosophical research to operationalize these requirements more precisely (see, e.g., [6, 5]). Encompassing some of these requirements, especially technical robustness, computational reliability came to be an important factor for trustworthiness in philosophical discussions (see, e.g., [Ferrario2022Explainability]).
@@
The last noteworthy discussion revolves around the relationship between explainable AI (XAI) and trust in AI. Increasing trust is one of—if not the—primary goal of XAI [Kaestner2021Relation, Langer2021What]. [Kaestner2021Relation] refer to this as the explainability-trust hypothesis [Kaestner2021Relation]. However, studies show that providing explanations does not consistently increase trust, and in many cases, even decreases it [Kaestner2021Relation, Langer2021What]. Here, trust calibration plays an important role, as explanations that reveal system errors should lead to lower trust. Unfortunately, psychological biases influence this process, as people sometimes trust a system simply because of the illusion of an explanation [DBLP:conf/chi/EibandBKH19]. On the same topic, [Ferrario2022Explainability] argue, drawing on the concept of anti-monitoring, that XAI cannot directly contribute to trust, as it constitutes a form of monitoring [Ferrario2022Explainability]. However, they argue that XAI can indirectly support trust because human-AI interaction that is supported by XAI is generally more trustworthy.
@@
Regarding biases and the possibility of generating problematic content, there is now more and more research that argues that, for a system to be adequately trusted, it is not just the system that needs to function correctly. Rather, for a system to be perceived as benevolent, there must be regulation, auditing processes, and established standards that govern the entire engineering process [13, manzini2024should, 8]. [13], for instance, argue that ethical governance, consisting of ethical codes of conduct, ethics training, responsible innovation, and transparency, is essential to build trust in AI systems [13]. Similarly, [manzini2024should] argue that user trust requires evidence on the functionality of an AI assistant at the level of its design, the practices of the developing organization, and oversight by external bodies [manzini2024should].
{results}
RQ1 (trust conceptualizations):
In psychology, the ability, benevolence, and integrity model of trust is common. How much (changes in) these factors affect trust depends on the trustor’s propensity to trust, which again affects the trustor’s initial level of trust. A central aspect is the risk perceived by a trustor when they trust the trustee to perform a certain task. The model was later extended to cover calibration, resolution, and specificity of trust. E.g., ill-calibrated trust may lead to disuse (in the case of undertrust) or misuse (in the case of overtrust) of a system.
In philosophy, anti-monitoring (not being able to monitor the trustee) is essential. Philosophers place more emphasis on the preconditions of trust and the conditions of trust formation than psychologists. They refer to calibrated trust as well-grounded trust and add the concepts justification and plausibility of trust. There are ongoing discussions whether trust can be meaningfully applied to AI.